OpenStudy (anonymous):

Let f(x)=x+5 and g(x)=x^2 find (g*f)(4) 9 81 20 1

OpenStudy (whpalmer4):

If \(f(x) = x+5\) and \(g(x) = x^2\) then \((g*f)(x)) = (x+5)^2\), right? Evaluate that at \(x=4\) and you've got your answer.

OpenStudy (anonymous):

So I would do (g*f)(4))=(4+5)^?

OpenStudy (anonymous):

(4+5)^2

OpenStudy (anonymous):

9^2=81

OpenStudy (anonymous):

so is 81 my answer?

OpenStudy (whpalmer4):

Yes, \[(g*f)(x) = g(f(x))\] \[(x+5)^2 = (4+5)^2 = 9^2 = 81\]

OpenStudy (anonymous):

Thank you so much!